# Accounting for Focus Ambiguity in Visual Questions

### v1

@@ 4 Model Benchmarking
We benchmark the modern model’s performance on VQ-Focus-Ambiguity for two novel tasks: (1) recognizing whether a visual question has focus ambiguity and (2) locating all image regions that could be the question’s focus.
@@ 4.1 Recognizing Questions with Focus Ambiguity
We benchmarked two models. First is InternVL2-Pro [7], the top-performing open-source model on the MMMU benchmark [41]. Second is GPT-4o [28], the top-performing model on the Arena [8] benchmark.
We used five prompts for each model. Three involved no supervision (i.e., zero-shot) while two incorporated a small amount of training data to ideally boost performance (i.e., in-context few-shot learning). They are as follows:
@@
  • Zero-shot chain of thought (ZS-CoT): augments ZS with the instruction, “please think step by step”.
  • Zero-shot enhanced chain of thought (ZS-ECoT): a prompt decomposition approach that augments ZS with structured guidance on completing the following four sequential steps: (1) understand the image, (2) understand the question, (3) find the target region(s) in the image, and (4) make a prediction. All details for enabling reproducibility are provided in the supplementary materials.
  • Few-shot (FS): augments ZS with two examples, one ambiguous and one unambiguous. Each example image is replaced with a textual description to maintain the benchmark consistency as some models don’t support multi-image input.
  • Few-shot enhanced chain of thought (FS-CoT): combines FS with the prompt decomposition of ZS (ECoT).
Because generative models can predict any arbitrary text output, we applied post-processing to convert all raw outputs to fit into three possible categories: “ambiguous”, “unambiguous”, and “undecided”. Post-processing details are provided in the supplementary materials.
@@
We employ five metrics. Four are standard binary classification metrics—accuracy, weighted average precision, weighted average recall, and weighted average f1 score— with the positive label being a visual question with focus ambiguity. Weighted scores mitigates biases that could arise from the slight class imbalance. Our fifth metric is “undecided rate”, which is the fraction of all examples with “undecided” predictions.
@@
Results are shown in Table 2.
Model Prompt Acc. P. R. F1 UR
ZS 31.6 22.0 55.7 21.0 0.0
ZS-CoT 33.0 33.6 33.1 32.8 0.0
GPT-4o ZS-ECoT 34.4 22.6 55.5 22.3 0.0
FS 35.0 24.4 59.4 22.5 0.0
FS-ECoT 36.4 23.8 56.7 23.5 0.3
ZS 51.3 32.0 64.8 31.2 1.6
ZS-CoT 52.1 32.8 65.3 31.8 1.8
InternVL2 ZS-ECoT 54.4 35.0 67.1 33.9 1.9
FS 54.0 35.0 67.7 34.6 1.0
FS-ECoT 53.6 34.2 66.6 33.3 1.4
Table 2: Performance of 10 model variants for question ambiguity recognition with respect to accuracy (Acc.), weighted average precision (P.), weighted average recall (R.), weighted average f1 score (F1), and undecided rate (UR) presented in percentage.
Figure 7: Fine-grained analysis on the performance of two benchmarked models with recognizing the presence of focus ambiguity for (a,b) each data source in VQA-FocusAmbiguity independently with respect to accuracy and F1-score respectively, (c) ambiguous versus unambiguous questions with respect to accuracy, (d) and answer length respect to accuracy. Also shown is the (e) percentage of predicted answers that are long versus short in the dataset.
Overall, all models perform poorly at question ambiguity recognition, with the top-performing model predicting the correct label only slightly more than half the time (i.e., for 54.4% of instances). This suggests our new dataset offers a challenging new problem for the research community.
When comparing across different prompt types, we observe that both models perform best overall (with respect to accuracy) when provided prompt decompositions demonstrating how to reason through the task. Specifically, the best performance is achieved with ZS-ECoT for InternVL2 and FS-ECoT for GPT-4o. We observe a consistent improvement in performance in ZS-CoT and ZS-ECoT compared to ZS, where both GPT-4o and InternVL2 demonstrate the performance order ZS-ECoT ¿ ZS-CoT ¿ ZS and Molmo ZS-CoT ¿ ZS-ECoT ¿ ZS. One valuable direction for further enhancing models is to dig into the step(s) of the reasoning process to establish which are the performance bottlenecks for such models.
When comparing performance across models, InternVL2 consistently outperforms GPT-4o. That is an exciting outcome as it means further advancements on today’s state-of-the-art can be achieved by any researcher, since InternVL2 is open-source rather than proprietary like GPT-4o.
@@
We next analyze the models’ performance with respect to (1) data source, (2) ambiguity label, and (3) answer length. Results are shown in Figure 7.
With respect to the data source, we observe that GPT-4o has higher accuracies on VQA-v2 and VizWiz-VQA, while InternVL2 has the best accuracies on MSRA-B (Figure 7(a)). In Figure 7(b), however, we see that the F1 scores are a lot more stable across data sources and prompts in InternVL2, while GPT-4o demonstrates large gaps between data source and prompts, suggesting that GPT-4o is relatively sensitive to data and prompt changes.
Whhn comparing models’ performance across ambiguous and unambiguous samples, we did not observe a significant trend for GPT-4o but found that InternVL2 performs consistently better on the unambiguous set than ambiguous set (Figure 7(c)). This reinforces the aforementioned analysis about data source, where we observed strong performance from InternVL2 on MSRA-B.
For answer length, short answers contained a single word, such as “ambiguous” or “unambiguous”, while long answers tended to provide extended reasoning. etter performance is typically achieved across all methods when answers are shorter (Figures 7(d) and (e)). Yet not always as, for both models, performance is better for long answers than for short answers in FS-ECoT. The tendency to generate long answers is especially pronounced across both models for ZS-CoT compared to other prompt types. From qualitative analysis, we found the models typically generate two to five sentences explaining their reasoning processes. Notably, the steps used in these answers are diverse and tailored to the specific questions and images. Some, but not all, resemble the steps we provide in ZS-CoT.
4.2 End-to-End FocLocating All Plausible Regions of Focus
@@
We evaluate two models. First is the state-of-the-art, end-to-end grounding model, GLaMM [32]. It is the top-performing model in Grounded Conversation Generation and Referring Expression Segmentation tasks, with top performance on the GranD and refCOCO benchmarks, respectively. We adopt the same prompting methods already defined in Section 4.1—ZS, ZS-CoT, ZS-ECoT, FS, and FS-ECoT—with the only modification that our problem definition indicates the task is to segment all relevant regions. We refer to this as the end-to-end approach.
We also benchmark an engineered solution that facilitates the reasoning process by breaking down the task into two simpler, sequential steps: describe and locate. We use the top-performing LLM, GPT-4o, for description generation and the state-of-the-art grounding model, GLaMM, for prompted grounding. We refer to this as the multi-step approach.
@@
We employ three metrics. First is the standard metric for instance segmentation: mAP. We also employ union IoU and max IoU to analyze models’ performance when they predict only one focus region. Specifically, we calculate union IoU as the intersection over union (IoU) between the predicted mask and the union of all focus regions to measure if the generated mask provides a semantic segmentation (rather than instance segmentations) that accurately captures all focus areas. We calculate max IoU as the largest mIoU score between the predicted mask and each focus region in the image, to see if the generated mask accurately captures a single focus region.
Figure 8: Fine-grained analysis on the performance of end-to-end and multi-step approaches for question grounding. (a) Union IoU scores of our four data sources. (b) Fine-grained analysis on data attributes: object and part in PACO and question-answer question grounding alignment. (c) An example of PACO parts grounded as focus regions in our dataset. (d) Union IoU scores of samples based on the number of focus regions.
@@
Table 3: Performance of two models for the focus ambiguity localization task with respect to three metrics: mAP for both approaches with the extension of union IoU and max IoU on the end-to-end approach.
Overall Results.
Results are shown in Table 3. The models perform consistently poorly in mAP across all prompting methods. We observe that while GLaMM supports multiple segmentation outputs, it consistently generates only a single mask for each sample in our test set. This limitation negatively impacts its end-to-end performance on samples with multiple grounding areas. While the multi-step approach has the potential to mitigate this gap, we observed that the number of descriptions generated by GPT-4o doesn’t have a strong correlation to the number of ground truth regions, so the effect is limited. We also observe a mild trend of max IoU being larger than union IoU in all tested prompts, suggesting that the model leans slightly toward looking for one of the focus regions in the prediction instead of the combination of all focus regions.
@@
We next analyze the performance for each data source independently as well as based on how many focus regions the visual questions contain in the ground truth. esults are shown in Figure 8.
We observe notable performance differences across dataset sources. Specifically, all models (end-to-end and multi-step GLaMM with five different prompt settings) perform best on MSRA-B, followed by VizWiz-VQA and VQA-v2, and lastly PACO-LVIS. Interestingly, this ranking is roughly the reverse of the GPT-4o’s performance in the focus ambiguity recognition task. We suspect that the high performance on MSRA-B is due to the ground truth often aligning with the most salient object, which aids localization. Our PACO-LVIS-based data often contain more focus regions with smaller areas, which can increase the grounding difficulty. These results highlight the potential of the integration between grounding and recognition models in future work.
Our results also show that IoU scores tend to drop in samples with a higher number of focus regions. Models perform significantly better on samples with a single grounding region than those with more regions, which aligns with InternVL2’s performance on ambiguity recognition.
Our fine-grained analysis of the models’ performance in locating PACO-LVIS’s objects versus parts shows that models struggle to identify parts across the methods, with considerably better performance observed for objects. This finding parallels progress in the broader computer vision community, where only relatively recently researchers have begun trying to segment parts. It also aligns with findings from an authentic use case [3], which shows that grounding models struggle most with small regions [3]—an attribute commonly associated with parts. Notably, the models perform consistently better on the samples where the question grounding aligns with the answer grounding.
@@ 5 Conclusion
We introduce a novel task with a dataset called VQ-Focus Ambiguity, which provides each region described in the question that is necessary to arrive at a valid answer. Our benchmarking results indicate that current models struggle to locate all plausible regions of focus when the focus is on parts, does not align with answer groundings, or when the number of foci is large. We share our dataset to facilitate future extensions of this work.
6 Supplementary Materials
@@ 7.1 Image Sources for PACO-LVIS
From preliminary analysis, we found that images with sufficient complexity for generating questions with focus ambiguity typically had either more than three segmented object instances or more than three segmented part instances. Therefore, our 1,700 samples sampled from the PACO-LVIS dataset were only those that met this complexity requirement.
@@ 7.2 Automated Candidate Question Generation
  • Default + Segmentation Overlay + Mirroring Requirements + Explanation: This prompt built on the previous setup by also asking the model to explicitly repeat the task requirements before generating questions and explain why the generated question satisfies the requirements.
@@ 8 Data Analysis
As discussed in the main paper, the answer grounding and question groundings can be different  [^6] We manually go through all 330 examples’ question and answer groundings to determine whether they match or differ. This is because all question groundings are annotated from scratch, and only considering IoU between question groundings and answer groundings might not provide an accurate evaluation. This is particularly true for small objects, where even slight boundary misalignments can significantly affect IoU scores. Additionally, question groundings and answer groundings may refer to the same object but exhibit boundary misalignments due to differences in annotators’ interpretations. In occluded scenarios, annotators might define regions differently—for example, one might include the occluded portion of an object, while another might exclude it—resulting in mismatches despite semantically correct predictions. . Additional examples are shown in Figure 15. As shown, the answer grounding and question groundings can be different because the question is asking about the relationship between things, (e.g., “What is the person standing on?” when the focus is “person” while the answer is “floor”/“carpet”) or locations of the entities (e.g., “Where is the vase placed?”). It also happens when the focus is clearly about an item but the answer can be features of the item (e.g., “What does this say?” when the focus is the product and the answer grounding is about the text label on the product). It also happens when there are multiple possible foci, but the answer only mentions the one that is fully visible compared to those partially visible; e.g., the one in the center compared to those not in the center; the one taking a larger fraction of the image compared to those that are smaller; the one in the foreground compared to those in the background; the one that is readable compared to those that are blurry.
@@
We provide the breakdown of the number of questions where question groundings and answer groundings match and are different with respect to the presence or absence of focus ambiguity (i.e., with multiple question groundings or single question grounding) and the presence or absence of answer ambiguity (i.e., with multiple answer groundings or single answer grounding) in Table 4. Qualitative examples are shown in Figure 16. This further supports the conclusions from Figure 15 that the differences between question groundings and answer groundings may arise when the question pertains to relationships between the entities, as shown in Figure 16 (d) (e) or locations of the entities, as shown in Figure 16 (c).
@@
Table 4: Number of questions that focus and answer groundings are matching and are different, respect to Single/Multiple Question Groundings and Single/Multiple Answer Groundings
9 Model Design
In model benchmarking, all models are tested with five prompting methods, zero-shot (ZS), zero-shot chain of thought (ZS-CoT), zero-shot enhanced chain of thought (ZS-ECoT), few-shot (FS), and few-shot enhanced chain of thought (FS-CoT). For focus ambiguity recognition, we prompted GPT-4o and InternVL2 using these five methods, respectively, to acquire classification results. For end-to-end region localization, we prompted GLaMM using these five methods to acquire segmentation masks. For multi-step localization, we prompted GPT-4o using these five methods to acquire descriptions of the regions, and then prompt GLaMM using the each generated description embedded in “Can you segment {description}?” to acquire segmentation masks. Below, we exemplify the detailed designs of these methods using the task, end-to-end localization.
Our introductory prompt starts every method by defining the task clearly and instructs the models only returning one word to simplify the post-processing of the classification results, which is: “You are a helpful assistant. You will be given an image and a question about the image. Your task is to predict whether the question is ambiguous or unambiguous based on the given image and the definition of focus ambiguity. Focus ambiguity in a visual question occurs when…(definition of focus ambiguity). Please only output “ambiguous” or “unambiguous”.”
The step-by-step instructions used in ZS-ECoT method are the following: “1. Understand the Image: Carefully analyze and fully comprehend the given image. 2. Understand the question: Carefully analyze and fully comprehend the question. 3. Find the regions: Find all regions in the image that could each satisfy all the constraints and contain all the necessary information to answer the question. 4. Predict: Predict whether the question has focus ambiguity or not. If the there are multiple regions found, it is an ambiguous question. Otherwise, it is an unambiguous question.”
In the few-shot methods, we provide two examples, one ambiguous and one unambiguous, in the prompts. The unambiguous example is: “[Question]: What color is the kite? [Image description]: The image shows a group of people in a park with picnic tables. One table has some food on it and there are two women sitting at the table. Other people are standing. One young boy is holding a blue kite in his hand, another child is picking up a red kite from the ground, and one man is running and flying a yellow kite. [Prediction]: ambiguous.” The ambiguous example is: “[Question]: Where is the man holding the apple? [Image description]: The image shows a man wearing a blue shirt sitting at a dining table and a little girl sitting on the floor beside him. The setting is likely a dining room. There is a window behind the dining table with brown curtains. The man is holding an apple in his hand. The little girl is holding a toy in her hand. A brown dog is laying on the floor looking at the girl. [Prediction]: unambiguous.”

### v2

@@ 4 Model Benchmarking
We now benchmark models on VQ-FocusAmbiguity for two novel tasks: (1) recognizing whether a visual question has focus ambiguity and (2) locating all image regions that could be the question’s focus.
@@ 4.1 Recognizing Questions with Focus Ambiguity
We evaluate four foundation models. Three are top-performers on Arena-Vision [10] and MMMU benchmarks [54]: GPT-4o [39], InternVL2-Llama3-76B [9], and Qwen2.5-VL-7B-Instruct [3]. The fourth is Molmo-7B [12], a state-of-the-art language grounding model.
Prompts.
We used five prompts for each model, resulting in 20 tested model variants. Three prompts involved no supervision (i.e., zero-shot) while two incorporated a small number of examples to ideally boost performance (i.e., in-context few-shot learning). They are as follows:
@@
  • Zero-shot chain of thought (ZS-CoT): facilitates the model’s reasoning by augmenting ZS with the instruction “please think step by step”.
  • Zero-shot enhanced chain of thought (ZS-ECoT): facilitates the model’s reasoning by augmenting ZS with structured guidance (i.e., four steps) on how to perform our novel task (i.e., prompt decomposition).
  • Few-shot (FS): augments ZS with an ambiguous and an unambiguous example. A textual description is used for each example image to maintain evaluation consistency, as some models don’t support multi-image input.
  • Few-shot enhanced chain of thought (FS-ECoT): augments FS with prompt decomposition from ZS-ECoT.
All prompts specify please only answer “ambiguous” or “unambiguous”, guiding the models to generate a one-word answer for binary classification. However, since generative models can and did produce arbitrary text beyond what was requested, we applied post-processing to categorize all outputs into three possible categories: “ambiguous”, “unambiguous”, and “undecided”.
@@
We employ four metrics. Two are standard binary classification metrics: accuracy and weighted f1 score. The third is “positive rate”, measuring the percentage of positive predictions (i.e., predicting there is question ambiguity) to reveal potential biases in model predictions. The last metric is “undecided rate”, which is the fraction of all examples with “undecided” predictions.
@@
Results are shown in Table 2. Overall, all models perform poorly, especially with respect to accuracy and F1 scores. This underscores that our dataset offers a challenging problem for the research community.
Model Prompt Acc. F1 Pos. UR
ZS 67.7 67.4 63.5 1.9
ZS-CoT 69.6 69.8 53.3 3.0
GPT-4o ZS-ECoT 68.4 68.6 46.8 2.8
(over 200B) FS 60.0 58.0 74.0 0.6
FS-ECoT 64.9 65.1 45.4 2.8
ZS 55.0 53.1 28.5 2.7
ZS-CoT 56.7 54.8 27.9 4.1
InternVL2 ZS-ECoT 54.9 53.2 29.6 3.1
(76B) FS 52.4 51.9 36.8 1.1
FS-ECoT 53.3 51.8 31.1 1.4
ZS 57.2 53.5 79.1 0.0
ZS-CoT 63.8 62.9 67.1 0.1
Qwen2.5-VL ZS-ECoT 65.5 65.3 59.0 0.0
(7B) FS 53.6 46.1 88.3 0.0
FS-ECoT 59.0 55.8 78.1 0.0
ZS 38.5 21.9 99.5 0.1
ZS-CoT 56.9 57.1 48.1 6.9
Molmo ZS-ECoT 45.7 42.6 75.0 1.0
(7B) FS 41.8 24.6 100.0 0.0
FS-ECoT 49.5 48.9 64.2 3.4
Table 2: Performance of 20 model variants for question ambiguity recognition with respect to accuracy (Acc.), weighted f1 score (F1), positive rate (Pos.), and undecided rate (UR) as percentages.
Figure 6: Fine-grained analysis of four models in recognizing focus ambiguity. Accuracy is reported based on data source and answer length. Percentages are provided relative to answer length. The five prompts are represented as follows: ZS (A), ZS-CoT (B), ZS-ECoT (C), FS (D), and FS-ECoT (E).
Our results also offer insights into strategies that can boost performance. First, we observe that facilitating models’ reasoning abilities through chain-of-thought prompts (i.e., CoT and ECoT methods) leads to considerable performance gains across all but one model (i.e., InternVL2). For instance, ZS-CoT prompting boosts Molmo-7B’s accuracy by 18.4 percentage points (pp), and ZS-ECoT yields an 8.3 pp improvement for Qwen2.5-VL. These gains elevate the performance of these 7B models to match or even surpass that of the much larger InternVL2 model (76B). We hypothesize the disparity observed for InternVL2 stems from the pretraining data, where both Qwen2.5-VL and Molmo were trained on region-level counting and pointing tasks included in the PixMo dataset [12], while InternVL2 was not trained. Such tasks are inherently related to ambiguity recognition: counting a single focus regions corresponds to an unambiguous question, and counting multiple focus regions indicate ambiguity. The positive rate scores provides evidence supporting our hypothesis, as InternVL2 consistently favors negative (i.e., unambiguous) predictions across prompting strategies while Molmo-7B and Qwen2.5-VL tend to classify questions as ambiguous under zero-shot conditions while adopting a more balanced perspective when guided by reasoning-based prompts. Together, these findings underscore the complementary importance of both the prompting strategy and training data in enhancing vision-language models’ ability to recognize question ambiguity.
@@
We next analyze the performance of the models with respect to data source and answer length. For data sources, we consider all image sources in VQA-FocusAmbiguity. For answer length, we categorize answers as short or long, where short answers contain one word (ideally “ambiguous” or “unambiguous”, as instructed in the prompt) and all other answers are long. Results are shown in Figure 6.
With respect to data source, we observe models exhibit similar performance on the three balanced datasets (i.e., PACO, VQA-v2, VizWiz-VQA) and different performance on the highly imbalanced MSRA-B dataset (i.e., only contains unambiguous examples). For instance, across all five tested prompts, InternVL2 consistently performs best on MSRA-B while GPT-4o and Qwen consistently performs worst on MSRA-B in comparison to the other three sources. Altogether, these findings underscore the models’ resilience to variations in image and question sources, while also highlighting they bring different prediction biases.
With respect to answer length, while we generally observe similar performance for short and long answers, a notable exception is for the best-performing model where performance for longer answers surpasses that of short answers when using chain-of-thought prompting (i.e., GPT-4o with the ZS-CoT prompt). This underscores that facilitating a model’s reasoning process can enhance overall performance. However, given our instruction to output a single word, a potential direction for future research to bridge the gap between model performance and user expectations could be to instead execute silent CoT reasoning to achieve better performance while still generating brief responses.
Figure 7: Molmo+SAM zero-shot results. Questions and ground truth masks are shown in Figures 1 and 5, stars denote where Molmo points, and blue masks denote SAM’s segmentations.
4.2 Locating All Plausible Regions of Focus
Figure 8: Fine-grained analysis on the performance of our three benchmarked approaches for question grounding using the five prompts, ZS (A), ZS-CoT (B), ZS-ECoT (C), FS (D), and FS-ECoT (E). (a) Union IoU scores of our four data sources. (b) Union IoU scores of object and part in PACO and the attribute based on question-answer question grounding alignment. (c) Union IoU scores of samples based on the number of focus regions. (d) An example of PACO parts grounded as focus regions in our dataset.
@@ 4.2 Locating All Plausible Regions of Focus
We evaluate three models. Included is the state-of-the-art language grounding model, GLaMM [43]. We also benchmark two engineered solutions that facilitate the reasoning process by breaking the task into two simpler, sequential steps. The first relies on GPT-4o to generate description of the focus regions and GLaMM to locate the regions given the descriptions. The second prompts Molmo [12] to generate points (i.e., x,y coordinates) locating all of a question’s focus regions and then feeds those as point prompts to SAM [28] to decode into segmentations.
Prompts.
For all models, we adopt the five prompting methods defined in Section 4.1, with minor modifications. For GLaMM, the problem definition instead indicates the task is to segment all focus regions. For ChatGPT-4o+GLaMM, we acquire descriptions of the focus regions from GPT-4o using the five prompts, and then acquire segmentations by passing each into a prompt for GLaMM specifying “Can you segment {description}?”. For Molmo+SAM, we replace the aforementioned prompt’s “segment” with “point”, since Molmo generates points.
@@
We employ three metrics. First is the standard metric for instance segmentation: mAP. We also employ union IoU and max IoU to analyze performance when models predict only one focus region. We calculate union IoU as the IoU between the predicted mask and the union of all focus regions to measure if the generated mask provides a semantic segmentation (rather than instance segmentations) capturing all focus regions. We calculate max IoU as the largest mIoU score between the predicted mask and each focus region, to see if the generated mask instead accurately captures a single focus region.
Overall Performance.
Results are shown in Table 3. All models perform poorly, with a considerable performance discrepancy across them.
@@
ZS 23.9 36.4 44.6
ZS-CoT 24.3 36.2 45.4
Molmo+SAM ZS-ECoT 24.2 36.1 44.3
FS 11.0 16.1 20.8
FS-ECoT - - -
Table 3: Performance of three models for focus ambiguity localization with respect to three metrics.
We attribute Molmo+SAM’s leading performance (in terms of mAP) to its tendency to point to multiple regions, as reinforced by Molmo’s nearly 100% positive rate observed for zero-shot settings in Table 2 indicating its bias to predict question ambiguity. Figure 7(b) shows where it successfully grounds two small clocks. However, Molmo+SAM features much lower union IoU, which is because SAM can fail to segment the whole object with only point input, such as for a “drinking duck” where only the leg of the duck is segmented rather than the whole duck. Additionally, Molmo encounters errors when processing FS-ECoT due to the long context input, which results in it not able to make predictions for FS-ECoT (denoted by “–”).
The other two models perform poorly for different reasons. While GLaMM supports multiple segmentation outputs, it consistently generates only a single mask. As for the GPT-4o-based model, its descriptions poorly correlated with the number of ground truth regions resulting in poor subsequent performance from GLaMM.
@@
We finally perform fine-grained analysis, with all results shown in Figure 8.
First, we observe notable performance differences across dataset sources (Figure 8a). Almost all models with different prompt settings perform best on MSRA-B, followed by VizWiz-VQA and VQA-v2, and lastly PACO. We attribute better performance on MSRA-B to the ground truth often aligning with the most salient object, which simplifies localization. In contrast, PACO-based data often contains many focus regions occupying smaller areas, which can increase the segmentation difficulty.
Our fine-grained analysis also show models struggle to identify parts, when comparing models’ performance in locating PACO’s objects versus parts (Figure 8(b)), as exemplified in Figure 8(d). This finding parallels progress in the broader computer vision community, where only relatively recently researchers have begun trying to segment parts.
Our results also underscore a correlation between performance and the number of focus regions, with worse performance when there are a greater number of focus regions (Figure 8(c)). This is exemplified in Figure 7(a).
Finally, we found models perform consistently worse on examples where the question grounding differ from answer grounding. This is exemplified in Figure 8(b) and Figures 7(c) and 7(d).
@@ 5 Conclusion
We introduced the VQ-FocusAmbiguity dataset for evaluating models’ abilities to acknowledge question ambiguity. Analysis reveals this dataset comes with unique challenges not explored in related VQA grounding datasets, particularly for answer grounding. Benchmarking reveals models struggle to recognize question ambiguity and locate all focus regions, underscoring the need for future research. We publicly-share the dataset to facilitate future progress.
@@
    User Interface to Mitigate Ambiguity (Supplements Section 1)
  2. 2.
@@
6 User Interface to Mitigate Ambiguity
This paper proposes a new directions for VQA by enabling the recognition of question ambiguity and localization of all focus regions. These can serve as valuable back-end capabilities to enable novel front-end human-computer interactions. For example, a front-end system could simply notify the user that ambiguity is detected, enabling the user to choose to try again with a modified question, image, or both. Alternatively, as exemplified in Figure 9, when ambiguity is detected, all plausible focus regions can be presented to the user for them to select one or more specific regions of interest through simple clicks.
Figure 9: Example of a possible user interaction interface that could help mitigate ambiguity with our focus ambiguity recognition and localization tasks. Specifically, when ambiguity is detected, the user interface will first present all plausible focus regions in the image and then prompt users to select one or more specific regions through simple clicks.
@@ 7.1 Image Sources for PACO-LVIS
From preliminary analysis, we found that images with sufficient complexity for generating questions with focus ambiguity typically had either more than three segmented object instances or more than three segmented part instances. Therefore, our 2,272 samples sampled from the PACO-LVIS dataset were only those that met this complexity requirement.
@@ 7.2 Automated Candidate Question Generation
  • Default + Segmentation Overlay + Mirroring Requirements + Explanation: This prompt built on the previous set-up by asking the model to also repeat the task requirements before generating questions and explain why the generated question satisfies the requirements.
@@
7.3 Annotation Rules Design
We designed the annotation rules as follows: (1) one author identified five rules  [^4] The rules addressed cases include single or multiple entities, demonstrative pronouns, entire-image references, and questions with options provided. by analyzing thousands of visual questions. Next, (2) three other authors independently annotated tens of diverse examples with these rules to identify all instances where their annotations differed, and then (3) the three authors refined the rules to prevent those differences going forward. The rules, provided alongside examples, clarified how they should be applied in practice by annotators during large-scale annotation collection.
@@ 8 Data Analysis
As discussed in the main paper, the answer grounding and question groundings can be different. To establish this, we manually reviewed all 330 examples’ question and answer groundings to determine whether they match or differ. This is because all question groundings are annotated from scratch, and only considering IoU between question groundings and answer groundings might not provide an accurate evaluation. This is particularly true for small objects, where even slight boundary misalignments can significantly affect IoU scores. Additionally, question groundings and answer groundings may refer to the same object but exhibit boundary misalignments due to differences in annotators’ interpretations. In occluded scenarios, annotators might define regions differently—for example, one might include the occluded portion of an object, while another might exclude it—resulting in mismatches despite semantically correct predictions.
Additional examples are shown in Figure 16. As shown, the answer grounding and question groundings can be different because the question is asking about the relationship between things, (e.g., “What is the person standing on?” when the focus is “person” while the answer is “floor”/“carpet”) or locations of the entities (e.g., “Where is the vase placed?”). It also happens when the focus is clearly about an item but the answer can be features of the item (e.g., “What does this say?” when the focus is the product and the answer grounding is about the text label on the product). It also happens when there are multiple possible foci, but the answer only mentions the one that is fully visible; e.g., the one in the center compared to those not in the center; the one taking a larger fraction of the image compared to those that are smaller; the one in the foreground compared to those in the background; the one that is readable compared to those that are blurry.
@@
We provide the breakdown of the number of questions where question groundings and answer groundings match and are different with respect to the presence or absence of focus ambiguity (i.e., with multiple question groundings or single question grounding) and the presence or absence of answer ambiguity (i.e., with multiple answer groundings or single answer grounding) in Table 4. Qualitative examples are shown in Figure 17. This further supports the conclusions from Figure 16 that the differences between question groundings and answer groundings may arise when the question pertains to relationships between the entities, as shown in Figure 17 (d) and (e) or locations of the entities, as shown in Figure 17 (c).
@@
Table 4: Number of questions where question and answer groundings are matching and different, with respect to Single/Multiple Question Groundings and Single/Multiple Answer Groundings.